Two filters described by the formulas below are cascaded.

Yı[n] = x[n] – x[n- 1] y2[n] = x[n] + x[n-1]

What will be the transfer function h(z) for the total system? from these four choices

H(Z) = 2 H(z) = 1 - 2-2 H(z) = 1 + 2.7-1 + 7-2 H(z) = 2-1
Answer #1

* Two filters are described by formeelas as 9.(n) = f(n) - X(n-1) 42(n) = x (h) + n(n-1) System / filter 1 :- (n) - (n) - (n-

42(2) = x(z) [1 +2 -] * Transfer function H₂(z) of filter 2 is ratio of output z- transforn 43(z) and input 7-transfour Xtz).